# Unveiling Factual Recall Behaviors of Large Language Models through Knowledge Neurons

### v2

@@ Abstract
Yifei Wang1,2 thanks: These authors contributed equally to this work., Yuheng Chen1,2 *, Wanting Wen1 , Yu Sheng1,2, Linjing Li1,2, Daniel Dajun Zeng1,2 1Institute of Automation, Chinese Academy of Sciences 2University of Chinese Academy of Sciences {wangyifei2022, chenyuheng2022}@ia.ac.cn {wanting.wen, shengyu2021, linjing.li, dajun.zeng}@ia.ac.cn
@@ 1 Introduction
Recent advancements in Large Language Models have underscored their exceptional reasoning prowess with natural language understanding across a broad spectrum of tasks Chen et al. (2023a); Kojima et al. (2022); Brown et al. (2020); Creswell et al. (2023). However, amidst these achievements, a specific form of reasoning has been somewhat overlooked and insufficiently investigated: reasoning tasks that entail the utilization of internal factual knowledge associations. For instance, when presented with a 2-hop question such as "Who is the chairperson of the manufacturer of the Holden Caprice?", LLMs must first identify the manufacturer of the Holden Caprice as General Motors, and subsequently retrieve the chairperson of General Motors from their internal knowledge or parametric knowledge Neeman et al. (2023); Zhong et al. (2024). Factual knowledge is observed to emerge in both GPT Meng et al. (2022) and Bert models Petroni et al. (2019); Jiang et al. (2020). Unlike mathematical Floyd (2007) and logical reasoning Pan et al. (2023), factual reasoning heavily relies on the factual knowledge encoded within LLMs, acquired through extensive pretraining on vast corpora, rather than on user-inputted premises. At the same time, it differs from commonsense reasoning Zhao et al. (2023); Trinh and Le (2019), which taps into general knowledge acquired through dynamic training to foster a holistic understanding of the world, instead of emphasizing specific factual information.
@@
Intuitively, it is reasonable to expect LLMs to harness their extensive internal knowledge to tackle reasoning tasks. Yet, an important question emerges: How effectively can LLMs actually retrieve and utilize their internal knowledge for reasoning purposes? Delving into this question is crucial for a multitude of reasons. First, LLMs’ efficient use of internal knowledge may greatly reduce reliance on external data sources, thereby lowering operational costs of data retrieval and API usage. Second, this dynamic capability allows the factual knowledge within LLMs to flow and interconnect Onoe et al. (2023), showcasing these models as organic entities rather than static information repositories Petroni et al. (2019). From a practical perspective, LLMs’ accurate retrieval and application of facts lead to more reliable and interpretable reasoning, enhancing their utility and trustworthiness in real-world applications.
@@
In this work, we investigate the harness of internal knowledge for reasoning through the lens of Knowledge Neurons (KNs). We focus on the basic setting of factual reasoning involving the composition of two facts (for example, "Who is the chairperson of the manufacturer of Holden Caprice?" in Figure 1). To achieve this, we carefully craft two-hop reasoning questions dataset that seamlessly integrates with the KN technique. We assess the level of factual recall at each reasoning step by introducing a novel metric, KN Scores. We examine KN Scores under three conditions of two-hop reasoning: no CoT, zero-shot CoT, and few-shot CoT, unveiling the pitfalls existing in the reasoning process and the enhancement effect of CoTWei et al. (2022). Then we conduct targeted interventions on KNs to enhance or suppress the factual retrieval process, finding the contributing impact on reasoning performance. Furthermore, we provide a detailed analysis of factual shortcuts Ju et al. (2024); Du et al. (2023), potentially caused by redundant information stored in models’ parameters within LLMs used for reasoning. Finally, we explore how the presence of knowledge conflict outside LLMs influences the factual recall process. Our findings can be summarized as follows:
@@ 2.2 Knowledge Neurons
Pretrained language models store vast amounts of factual knowledge and have a strong ability to recall this factual knowledge without further training Petroni et al. (2019); Jiang et al. (2020). Drawing inspiration from the key-value-memory nature of feed-forward layers Geva et al. (2021), Dai et al. (2022) proposes that factual knowledge is stored in specific neurons within the Feed-Forward Networks (FFNs) of the Transformer models, termed as knowledge neurons. They find that knowledge neurons are activated by knowledge-expressing prompts. The higher the activation of these knowledge neurons is, the more significantly their corresponding facts are expressed. Therefore, to assess the recall and utilization of the fact triplet $(s,r,o)$ necessary in the reasoning process, we refer to the activity of KNs as an indicator of factual recall. We make the following invariant assumptions: the KNs responsible for the expression of particular relational facts remain consistent across different application contexts. A specific fact is indicated by the same set of KNs under both single-hop queries and reasoning queries, which is a cornerstone for subsequent experiments. In Appendix B, We detail a methodology utilizing integrated gradients Sundararajan et al. (2017) to compute the contribution of all neurons in the intermediate layers of FFNs to the correct prediction of a multi-token ground truth, identifying neurons with greater contributions as KNs.
@@ 3 TFRKN: Two-hop Factual Reasoning for Knowledge Neurons
To investigate the behavior of factual recall in reasoning tasks for LLMs, we have developed a dataset called TFRKN (Two-hop Factual Reasoning for Knowledge Neurons). This dataset is composed of two-hop factual questions, which are constructed with frequently occurring entities Mallen et al. (2023) in Wikidata Vrandečić and Krötzsch (2014) and manually selected relations. The construction method is detailed in Appendix A. TFRKN dataset encompasses 4,550 distinct instances that cover 213 unique relational combinations. In alignment with the KN methodology, we have reformulated each fact triplet into over five varied query forms with the aim of refining true-positive KNs from specific-form queries. (An instance in TFRKN is shown in Table 6).
@@ 4.1 KN Scores
To quantify the capacity for internal recall of specific facts within LLMs, we devise a novel metric, termed KN Scores, as follows:
@@
where $\omega_{i}^{l}$ denotes the $i$-th neuron in the $l$-th intermediate layer of FFN and symbol $\omega$ represents the KNs associated with a specific fact triplet, denoted as $(s,r,o)$. For the first-hop fact and the second-hop fact, we designate their respective sets of KNs as $\omega_{1}$ and $\omega_{2}$. Under the context of a single-hop query, we denote KN Scores as $\{\overline{\omega}|QT_{1H}\}$. Similarly, within the two-hop reasoning context, KN Scores are represented as $\{\overline{\omega}|QT_{2H}\}$.
Figure 2: Scaled visualization of neuron activity within the intermediate layers of FFNs in Mistral-7B for the same case (A 32-layer×$\times$14336-neuron matrix). The vertical axis shows layer depth, and the horizontal axis shows neuron index in the FFN’s intermediate layers.
@@ 4.2 Experiment
We begin by filtering out reasoning questions where LLMs are unable to recall all individual facts, ensuring that any reasoning failures are due to the models’ inability to retrieve factual information rather than a lack of the foundational knowledge necessary for performing reasoning tasks. We then proceed to employ $\text{Fact}_{1}\text{Query}$ and $\text{Fact}_{2}\text{Query}$ (in Table 6) from each data point to pinpoint the positions of KNs for each-hop fact. Then we hook the values of each neuron belonging to $\omega_{1}$ and $\omega_{2}$ across various query scenarios to compute KN Scores. Using the KN Scores metric, we evaluate the recall of each fact under three distinct experimental conditions: no CoT, zero-shot CoT, and few-shot CoT. For each condition, we record KN Scores for both the first-hop $\{\overline{\omega}_{1}|QT_{2H}\}$ and the second-hop $\{\overline{\omega}_{2}|QT_{2H}\}$ facts within the context of two-hop reasoning questions. We select the KN Scores $\{\overline{\omega}_{1}|QT_{1H}\}$ and $\{\overline{\omega}_{2}|QT_{1H}\}$ under single-hop queries as baselines since KNs are significantly active in that straightforward context. We experiment with the instructed versions of three popular open-source models: LLaMA2-7B Touvron et al. (2023), LLaMA3-8B, Mistral-7B Jiang et al. (2023) (see Appendix D for more experimental details).
@@ 4.3 Results
Table 1: KN Scores for three conditions across three models. $\overline{\omega}$ is the KN Score of a particular fact while $\Delta$ indicates the change ratio (in percentages) of values compared with the single-hop baselines.
@@
CoT, whether zero-shot or few-shot, markedly improves factual knowledge utilization in LLMs over no CoT (see a case in Figure 2), which is evidenced by a higher $\Delta_{\overline{\omega}_{1}}$ and $\Delta_{\overline{\omega}_{2}}$ compared with no CoT setting, as shown in Table 1. We posit that this enhancement is likely driven by the step-by-step thinking process, which further stimulates the recall of facts as multi-hop reasoning progresses. This hypothesis can be supported by comparing the zero-shot CoT and few-shot CoT settings. Across three models, it is clear that zero-shot CoT struggles to significantly improve the recall of the second-hop fact compared to the reinforcement of the first-hop fact recall. However, consistent improvement across both triplets can be observed for few-shot settings. This observation strongly suggests that the reasoning direction in zero-shot scenarios is unclear, which prevents models from effectively identifying which relations of facts concerning the bridge entity to retrieve. In stark contrast, few-shot scenarios often mitigate this issue. Through the acquisition of knowledge from contextual demonstrations, models are more inclined to determine the subsequent phase in the reasoning trajectory and, in turn, adeptly utilize the relevant factual information via their attention mechanisms.
@@
Figure 3 illustrates a positive correlation between the recall of relevant fact triplets and reasoning accuracy. This relationship is especially pronounced in the case of LLaMA3-8B model under few-shot CoT, where the maximum increase in the recall of both $\Delta_{\overline{\omega}_{1}}$ and $\Delta_{\overline{\omega}_{2}}$ leads to the highest reasoning accuracy. The eliciting effect of CoT on factual recall across various LLMs is not uniform. For instance, zero-shot CoT mitigates the forgetting of factual information to some extent for LLaMA2-7B, whereas for LLaMA3-8B, zero-shot CoT enhances the retrieval of factual information to a level comparable to few-shot CoT. This adequately illustrates that the efficacy of CoT is also contingent upon the intrinsic capabilities of the LLMs themselves when they are of nearly the same scale.
@@ 5.2 Experiment
$\operatorname{ER}=\frac{\mathbf{|}\{\zeta\mid\operatorname{F}_{{\theta^{\prime}}}(\operatorname{QT}_{r_{2}}(r_{1}(s))=o_{2}\}\mathbf{|}}{|\Omega_{F}|},\forall\zeta\in\Omega_{F}$
$\operatorname{SR}=\frac{\mathbf{|}\{\zeta\mid\operatorname{F}_{{\theta^{\prime\prime}}}(\operatorname{QT}_{r_{2}}(r_{1}(s))\neq o_{2}\}\mathbf{|}}{|\Omega_{T}|},\forall\zeta\in\Omega_{T}$
@@ 5.3 Results
Ehan. $\Delta$ ER $\Delta$ ER $\Delta$ ER
@@ 6.1 Experiment
6.2 Results Analysis
Figure 4 provides a closer look at the shortcut phenomenon. The percentage of FF is significantly low, illustrating that it is hard for LLMs to fail to retrieve any factual information relevant when presented with the clues of overlapping entities or relational vocabulary in queries. For most instances of shortcuts, LLMs prefer to utilize the second-hop fact to directly answer reasoning questions, skipping the intermediate reasoning steps and relying on the object $o_{2}$ in the second-hop to cheat (a high ratio for FT). For TF cases, there might exist the direct associations between the head entity $s$ and the tail entity $o_{2}$ leveraged to derive correct answers.
7 Impact of Contextual Conflict
7.1 Experiment
Models Mistral-7B LLaMA2-7B LLaMA3-8B
MH SC MH SC MH SC
No CoT 55.75 44.25 60.51 39.49 71.89 28.11
Zero-shot 70.84 29.16 64.26 35.74 95.66 4.34
Few-shot 91.23 8.77 89.02 10.98 97.65 2.35
Table 5: Results of taking shortcuts. MH denotes successful retrieval of both facts while others denote by SC.
Setup
For each data point, we formulate a single-hop conflict fact by devising a set of potential objects denoted as $O_{candi}$ for its $r$. From this set, we deliberately select an object $o^{*}\neq o$ to introduce a knowledge conflict. In contradistinction, we also fabricate an entirely unrelated fact for each data point to serve as a distractor, referred to as knowledge distraction. We then respectively append the knowledge conflict and knowledge distraction sentences before the two-hop question, which is input into LLMs. Then we observe the values of KN Scores for each-hop fact. The examples of knowledge conflict and distraction for the first-hop and the second-hop facts are shown in Table 4.

### v3

@@ Abstract
Yifei Wangthanks: Equal contribution.1,2 ,Yuheng Chen11footnotemark: 12 ,Wanting Wen1 ,Yu Sheng1,2,Linjing Li thanks: Corresponding author.1,2,3, Daniel Zeng1,2 1 State Key Laboratory of Multimodal Artificial Intelligence Systems, Institute of Automation, Chinese Academy of Sciences, Beijing, China 2 School of Artificial Intelligence, University of Chinese Academy of Sciences, Beijing, China 3 Beijing Wenge Technology Co., Ltd, Beijing, China {wangyifei2022, chenyuheng2022}@ia.ac.cn {wanting.wen, shengyu2021, linjing.li, dajun.zeng}@ia.ac.cn
@@ 1 Introduction
Recent advancements in Large Language Models have underscored their exceptional reasoning prowess with natural language understanding across a broad spectrum of tasks Chen et al. (2023a); Kojima et al. (2022); Brown et al. (2020); Creswell et al. (2023). However, amidst these achievements, a specific form of reasoning has been somewhat overlooked and insufficiently investigated: reasoning tasks that require the utilization of internal factual knowledge associations. For instance, when presented with a 2-hop question such as "Who is the chairperson of the manufacturer of the Holden Caprice?" in Figure 1, LLMs must first identify that the manufacturer of the Holden Caprice is General Motors, and subsequently retrieve the name of General Motors’ chairperson from their internal knowledge, also referred to as parametric knowledge Neeman et al. (2023); Zhong et al. (2024). Previous work has shown factual knowledge emerges in both GPT Meng et al. (2022) and Bert models Petroni et al. (2019); Jiang et al. (2020). Unlike mathematical Floyd (2007) and logical reasoning Pan et al. (2023), factual reasoning heavily relies on the factual knowledge encoded within LLMs, acquired through extensive pretraining on vast corpora, rather than on user-inputted premises. At the same time, it differs from commonsense reasoning Zhao et al. (2023); Trinh and Le (2019), which taps into general knowledge acquired through dynamic training to foster a holistic understanding of the world, instead of emphasizing specific factual information.
@@
Intuitively, it is reasonable to expect LLMs to harness extensive parametric knowledge to tackle reasoning tasks. Yet, an important question emerges: How effectively can LLMs actually retrieve and utilize their internal knowledge for reasoning purposes? Delving into this question is crucial for several reasons. First, efficient use of parametric knowledge may significantly reduce reliance on external data sources, thereby lowering operational costs of data retrieval and API usage. Second, this dynamic capability allows the knowledge within LLMs to flow and interconnect Onoe et al. (2023), showcasing these models as organic entities rather than static information repositories Petroni et al. (2019). From a practical perspective, the accurate retrieval and application of parametric knowledge lead to more reliable and interpretable reasoning, enhancing their utility and trustworthiness in real-world applications.
@@
In this work, we investigate the harness of internal knowledge for reasoning through the lens of Knowledge Neurons (KNs). We focus on the basic setting of factual reasoning involving the composition of two facts (for example, "Who is the chairperson of the manufacturer of Holden Caprice?" in Figure 1). To achieve this, we carefully craft two-hop reasoning questions dataset that seamlessly integrates with the KN technique. We assess the level of factual recall at each reasoning step by introducing a novel metric, KN Scores. We examine KN Scores under three conditions of two-hop reasoning: no CoT, zero-shot CoT, and few-shot CoT, unveiling the pitfalls existing in the reasoning process and the enhancement effect of CoT Wei et al. (2022). Then we conduct targeted interventions on KNs to enhance or suppress the factual retrieval process, finding the contributing impact on reasoning performance. Furthermore, we provide a detailed analysis of factual shortcuts Ju et al. (2024); Du et al. (2023); Li et al. (2024), potentially caused by redundant information stored in models’ parameters within LLMs used for reasoning. Finally, we explore how the presence of knowledge conflict outside LLMs influences the factual recall process. Our findings can be summarized as follows:
@@ 2.2 Knowledge Neurons
Pretrained language models store vast amounts of factual knowledge and have a strong ability to recall this factual knowledge without further training Petroni et al. (2019); Jiang et al. (2020). Drawing inspiration from the key-value-memory nature of feed-forward layers Geva et al. (2021), Dai et al. (2022) proposes that factual knowledge is stored in specific neurons within the Feed-Forward Networks (FFNs) of the Transformer models, termed as knowledge neurons. They find that knowledge neurons are activated by knowledge-expressing prompts. The higher the activation of these knowledge neurons is, the more significantly their corresponding facts are expressed. Therefore, to assess the recall and utilization of the fact triplet $(s,r,o)$ necessary in the reasoning process, we refer to the activity of KNs as an indicator of factual recall. We make the following invariant assumptions: the KNs responsible for the expression of particular relational facts remain consistent across different application contexts. A specific fact is indicated by the same set of KNs under both single-hop queries and reasoning queries, which is a cornerstone for subsequent experiments. In Appendix B, We detail a methodology that utilizes integrated gradient Sundararajan et al. (2017) method to compute the contribution of all neurons in the intermediate layers of FFNs to the correct prediction of a multi-token ground truth, identifying neurons with greater contributions as KNs.
@@ 3 TFRKN: Two-hop Factual Reasoning for Knowledge Neurons
To investigate the behavior of factual recall in reasoning tasks for LLMs, we have developed a specialized dataset for knowledge neurons called Two-hop Factual Reasoning for Knowledge Neurons, TFRKN.
Dataset Construction
Our dataset consists of two-hop factual questions, where each question involves two facts that are connected by an intermediate entity. LLMs are more likely to recall triplets related to popular entitiesMallen et al. (2023). Therefore, for entity selection, we use the cumulative pageview count over the past 12 months as a metric and select the top 500 popular entities from Wikidata Vrandečić and Krötzsch (2014) based on this criterion. Two-hop fact triplets are then extracted from sub-graphs consisting solely of a set of manually selected relations and entities. To identify KNs for each-hop fact, we reformulate each fact triplet into more than five varied natural questions using ChatGPT (Appendix A). The TFRKN dataset encompasses 4,550 distinct instances covering 213 unique relational combinations with a sample instance shown in Table 6.
@@ 4.1 KN Scores
To evaluate the efficacy of factual recall within LLMs during reasoning tasks, we devise a novel metric termed KN Scores as follows:
@@
where $\operatorname{H}^{(l)}$ represents the input to the FFN of the $l$-th layer, which consists of the outputs from the $l$-th attention layer combined with the residual stream; $\omega_{i}^{l}$ denotes the $i$-th neuron in the $l$-th intermediate layer of FFN; $\omega$ represents the KNs associated with a specific fact triplet, denoted as $(s,r,o)$; $|\omega|$ denotes the size of the set, i.e., the number of KNs; and $\operatorname{SiLU}$ denotes the activation function. For the first-hop and second-hop fact, we designate their respective sets of KNs as $\omega_{1}$ and $\omega_{2}$. Under the context of a single-hop query, we denote KN Scores as $\{\overline{\omega}|QT_{1H}\}$. Similarly, within the two-hop reasoning context, KN Scores are represented as $\{\overline{\omega}|QT_{2H}\}$.
Figure 2: Scaled visualization of neuron activities within the intermediate layers of FFNs in Mistral-7B for the same case (A 32-layer×$\times$14336-neuron matrix). The vertical axis shows the depth of layers, while the horizontal axis shows the neuron index in the FFN’s intermediate layers. It is evident that KNs are distributed in the middle and final layers.
@@ 4.2 Experiment
We begin by filtering out reasoning questions where LLMs are unable to recall all individual facts, ensuring that any reasoning failures are due to the models’ inability to retrieve factual information rather than a lack of the foundational knowledge necessary for performing reasoning tasks. We then proceed to employ $\text{Fact}_{1}\text{Query}$ and $\text{Fact}_{2}\text{Query}$ (in Table 6) from each data point to pinpoint the positions of KNs for each-hop fact. Then we hook the values of each neuron belonging to $\omega_{1}$ and $\omega_{2}$ across various query scenarios to compute KN Scores. Using the KN Scores metric, we evaluate the recall of each fact under three distinct experimental conditions: no CoT, zero-shot CoT, and few-shot CoT. For each condition, we record KN Scores for both the first-hop $\{\overline{\omega}_{1}|QT_{2H}\}$ and the second-hop $\{\overline{\omega}_{2}|QT_{2H}\}$ facts within the context of two-hop reasoning questions. We select the KN Scores $\{\overline{\omega}_{1}|QT_{1H}\}$ and $\{\overline{\omega}_{2}|QT_{1H}\}$ under single-hop queries as baselines since KNs are significantly active in that straightforward context. We experiment with the instructed versions of three popular open-source models: LLaMA2-7B Touvron et al. (2023), LLaMA3-8B, Mistral-7B Jiang et al. (2023) (see Appendix C for more experimental details).
@@ 4.3 Results
Table 1: KN Scores for three conditions across three models. $\overline{\omega}$ is the KN Score of a specific fact while $\Delta$ indicates the change ratio (in percentages) of values compared with the single-hop baselines.
@@
CoT, whether zero-shot or few-shot, markedly improves factual knowledge utilization in LLMs over no CoT (KNs are more activated under CoT settings in Figure 2), which is evidenced by a higher $\Delta_{\overline{\omega}_{1}}$ and $\Delta_{\overline{\omega}_{2}}$ compared with no CoT setting, as shown in Table 1. We posit that this enhancement is likely driven by the step-by-step thinking process, which further stimulates the recall of facts as multi-hop reasoning progresses. This hypothesis can be supported by comparing the zero-shot and few-shot CoT settings. Across three models, it is clear that zero-shot CoT struggles to significantly improve the recall of the second-hop fact compared to the reinforcement of the first-hop fact recall. However, consistent improvement across both triplets can be observed for few-shot settings. This observation strongly suggests that the reasoning direction in zero-shot scenarios is unclear, which prevents models from effectively identifying which relations of facts concerning the bridge entity to retrieve. In stark contrast, few-shot scenarios often mitigate this issue. Through the acquisition of knowledge from contextual demonstrations, models are more inclined to determine the subsequent phase in the reasoning trajectory and, in turn, adeptly utilize the relevant factual information via their attention mechanisms.
@@
The combination of Figure 3 and Table 1 illustrates a positive correlation between the recall of relevant fact triplets and reasoning accuracy. This relationship is especially pronounced in the case of LLaMA3-8B model under few-shot CoT, where the maximum increase in the recall of both $\Delta_{\overline{\omega}_{1}}$ and $\Delta_{\overline{\omega}_{2}}$ leads to the highest reasoning accuracy. However, the eliciting effect of CoT on factual recall across various LLMs is not uniform. For instance, zero-shot CoT mitigates the forgetting of factual information to some extent for LLaMA2-7B, whereas for LLaMA3-8B, zero-shot CoT enhances the retrieval of factual information to a level comparable to few-shot CoT. This adequately illustrates that the efficacy of CoT is also contingent upon the intrinsic capabilities of the LLMs themselves.
@@ 5.2 Experiment
$\operatorname{ER}=\frac{\mathbf{|}\{\zeta\mid\operatorname{F}_{{\theta^{\prime}}}(\operatorname{QT}_{r_{2}}(r_{1}(s)))=o_{2}\}\mathbf{|}}{|\Omega_{F}|},\forall\zeta\in\Omega_{F}$
$\operatorname{SR}=\frac{\mathbf{|}\{\zeta\mid\operatorname{F}_{{\theta^{\prime\prime}}}(\operatorname{QT}_{r_{2}}(r_{1}(s)))\neq o_{2}\}\mathbf{|}}{|\Omega_{T}|},\forall\zeta\in\Omega_{T}$
@@ 5.3 Results
Enha. $\Delta$ ER $\Delta$ ER $\Delta$ ER
@@ 6 Analysis of Shortcuts
6.1 Experiment
Setup
Models Mistral-7B LLaMA2-7B LLaMA3-8B
MH SC MH SC MH SC
No CoT 55.75 44.25 60.51 39.49 71.89 28.11
Zero-shot 70.84 29.16 64.26 35.74 95.66 4.34
Few-shot 91.23 8.77 89.02 10.98 97.65 2.35
Table 4: The fraction of MH and SC in correctly answered examples (TT+FT). MH denotes successful retrieval of both facts while others denote by SC.
6.2 Results Analysis
Figure 4 provides a closer look at the shortcut phenomenon. The percentage of FF is significantly low, illustrating that it is hard for LLMs to fail to retrieve any factual information relevant when presented with the clues of overlapping entities or relational vocabulary in queries. For most instances of shortcuts, LLMs prefer to utilize the second-hop fact to directly answer reasoning questions, skipping the intermediate reasoning steps and relying on the object $o_{2}$ in the second-hop to cheat (a high ratio for FT). For TF cases, there might exist direct associations between the head entity $s$ and the tail entity $o_{2}$ leveraged to derive correct answers. In conclusion, experimental results show that recalling two-hop facts (TT) benefits the model’s reasoning performance. Specifically, with the presence of CoT, the proportion of TT significantly increases and the model’s reasoning accuracy improves substantially.
7 Impact of Contextual Conflict
@@ 7.1 Experiment
For each data point, we formulate a single-hop conflict fact by devising a set of potential objects denoted as $O_{candi}$ for its $r$. From this set, we deliberately select an object $o^{*}\neq o$ to introduce a knowledge conflict. In contradistinction, we also fabricate an entirely unrelated fact for each data point to serve as a distractor, referred to as knowledge distraction (See detailed construction in Appendix D). We then respectively append the knowledge conflict and knowledge distraction sentences before the two-hop question under no CoT setting, which is input into LLMs. Then we observe the values of KN Scores for each-hop fact. The examples of knowledge conflict and distraction for the first-hop and the second-hop facts are shown in Table 5.